### v4

# Thermodynamics of black and white holes in ensemble of Planckons

Abstract
The Tsallis-Cirto non-extensive statistics with $\delta=2$ describes the processes of splitting and merging of black holes and their thermodynamics.Volovik2025a ; Volovik2025d Here we consider a toy model, which matches this generalized statistics and extends it by providing the integer valued entropy of the black hole, $S_{\rm BH}(N)=N(N-1)/2$. In this model the black hole consists of $N$ the so-called Planckons – objects with reduced Planck mass $m_{\rm P}=1/\sqrt{8\pi G}$ – so that its mass is quantized, $M=Nm_{\rm P}$. The entropy of each Planckon is zero, but the entropy of black hole with $N$ Planckons is provided by the $N(N-1)/2$ degrees of freedom – the correlations between the gravitationally attracted Planckons. This toy model can be extended to a charged Reissner-Nordström (RN) black hole, which consists of charged Planckons. Despite the charge, the statistical ensemble of Planckons remains the same, and the RN black hole with $N$ Planckons has the same entropy as the electrically neutral hole, $S_{\rm RNBH}(N)=N(N-1)/2$. This is supported by the adiabatic process of transformation from the RN to Schwarzschild black hole by varying the fine structure constant. The adiabaticity is violated in the extreme limit, when the gravitational interaction between two Planckons is compensated by the repulsion between their electric charges, and the RN black hole loses stability. The entropy of a white hole formed by the same $N$ Planckons has negative entropy, $S_{\rm WH}(N)=-N(N-1)/2$.
pacs:
I Introduction
In references Volovik2025a ; Volovik2025d we considered the black hole thermodynamics using the non-extensive statistics. The entropy of the black hole obeys the composition rule
$\sqrt{S_{\rm BH}(M_{1}+M_{2})}=\sqrt{S_{\rm BH}(M_{1})}+\sqrt{S_{\rm BH}(M_{2})}\,,$ (1)
which coincides with the composition rule in the non-extensive Tsallis-Cirto $\delta=2$ statistics.TsallisCirto2013 ; Tsallis2020 Eq.(1) expresses the entropy of the black hole with mass $M_{1}+M_{2}=M$, which is obtained by the merging of two black holes with masses $M_{1}$ and $M_{2}$. In the process of merging the entropy increases. The reversed process of splitting of the black hole with mass $M$ into two smaller black holes with $M_{1}+M_{2}=M$ has been calculated using the quantum tunneling approach developed by Parikh and Wilczek.Wilczek2000 These calculations showed that the rate of this process can be expressed in terms of the difference in entropy before and after splitting, $w\propto\exp(-\Delta S)$, which confirms the composition rule in the equation (1). Then, following the Landau-Lifshitz book ”Statistical Physics”Landau_Lifshitz , one can assume that the splitting of a black hole in the process of macroscopic quantum tunneling is equivalent to a thermodynamic fluctuation in which the entropy decreases after splitting and then is restored by merging.
Here we consider a toy model, which gives rise to quantization of the generalized black hole entropy. The black hole is considered as an ensemble of the Planck mass objects – Planckons.Markov1967 ; Hawking1971 ; Aharonov1987 ; Rovelli2024 ; YenChinOng2025 ; Trivedi2025 If Planckon has the reduced Planck mass, $m_{\rm P}=1/\sqrt{8\pi G}$, the black hole with mass $M=Nm_{\rm P}$ consisting of $N$ Planckons has the integer valued entropy $S_{\rm BH}(N)=N(N-1)/2$. This leads to the extension of the composition rules, in which the processes of merging and splitting of black holes are expressed in terms of the integer numbers. For example, the probability of splitting of a black hole into two smaller parts, $N=N_{1}+N_{2}$, is given by the exponent $\exp(-\Delta S)=e^{-N_{1}N_{2}}$. For $N_{1}=1$ and $N_{2}=N-1$ this gives the rate of emission of single Planckons. This rate automatically includes the effect of back reaction in the process of radiation, which was first considered by Parikh and Wilczek.Wilczek2000
We also consider the extension of the generalized entropy to white holes and to the Reissner-Nordström (RN) black holes. We find that the statistical ensemble of charged $N$ Planckons forming the RN black hole has the same entropy as the electrically neutral black hole, $S_{\rm RNBH}(N)=S_{\rm BH}(N)=N(N-1)/2$. This is valid up to the extreme limit, after which the RN black hole loses stability. For a white hole, the anti-symmetry with respect to time reversal gives negative entropy, $S_{\rm WH}(N)=-S_{\rm BH}(N)=-N(N-1)/2$.
II Schwarzschild black hole
II.1 Horizon, irreversibility and thermodynamics
We use the Painleve-Gullstrand coordinates,Painleve ; Gullstrand where the metric has the following form:
$ds^{2}=-dt^{2}(1-{\bf v}^{2})-2dt\,d{\bf r}\cdot{\bf v}+d{\bf r}^{2}\,.$ (2)
Here the vector ${\bf v}$ is the shift function in the Arnowitt-Deser-Misner formalism.ADM2008 , while in the condensed matter analogs this vector corresponds to the ”velocity of the quantum vacuum”. For the Schwarzschild black and white holes the shift function is:
${\bf v}({\bf r})=\mp\hat{\bf r}\sqrt{\frac{2MG}{r}}\,.$ (3)
The minus sign in the equation (3) corresponds to the black hole configuration (the velocity ${\bf v}$ is directed toward the horizon), and the plus sign describes a white hole (the ”vacuum” flows away from the horizon). In both cases the event horizon is at $r=2MG$.
It is the event horizon which is responsible for irreversibility and thus for the thermodynamic behavior of the black and white holes.Witten2025 The external particle is absorbed by the black hole with probability 1, while the reversed process – the emission from the black hole – is suppressed by the Hawking exponent describing the rate of the quantum tunneling from the black hole. The detailed balance gives rise to the Hawking temperature of the black hole and to the Hawking temperature with minus sign for the white hole.
On the other hand the processes of merging and splitting of the black holes by macroscopic quantum tunneling determine the entropy of these objects. They obey the Tsallis-Cirto statistics, where the entropy is proportional to the square of the logarithm of probability:
$S_{\delta=2}=\sum_{i}p_{i}\left(\ln\frac{1}{p_{i}}\right)^{2}\,.$ (4)
II.2 Entropy in terms of Planck-scale elements
This peculiar type of the Tsallis-Cirto statistics (i.e. with $\delta=2$ in Eq.(4)) also suggests that the black hole can be represented as ensemble of $N$ ”black hole atoms” – the smallest collapsed objects introduced by Hawking,Hawking1971 or ”Planckons”,Aharonov1987 ; YenChinOng2025 see also Markov’s maximons.Markov1967 Planckon is the object with the Planck scale mass. The proper choice of the Planckon mass, which naturally fits the the Tsallis-Cirto statistics, is the reduced Planck mass, $m_{\rm P}^{2}=1/8\pi G$, so that the mass of the black hole is
$M=Nm_{\rm P}\,\,,\,\,m_{\rm P}=1/\sqrt{8\pi G}\,,$ (5)
This allows the Tsallis-Cirto entropy to be expressed in terms of an integer number of Planckons:
$S_{\rm BH}(N)=\begin{pmatrix}N\\ 2\end{pmatrix}=\frac{N!}{2!(N-2)!}=\frac{N(N-1)}{2}\,\,,\,\,N=\frac{M}{m_{\rm P}}\,.$ (6)
This entropy coincides with the Bekenstein-Hawking entropy in the limit of large $N$. On the other hand, it also describes the entropy of an ensemble with finite $N$.
In Eq.(6), the entropy of the ensemble arises from the correlations between Plancons. It can be interpreted as the number $\cal N$ of ways to select two Planckons from an ensemble of $N$ Planckons:
$S_{\rm BH}(N)={\cal N}=\frac{N(N-1)}{2}\,.$ (7)
In this interpretation, $N$ Planckons interact gravitationally with each other, and their correlated (or entangled) pairs provide ${\cal N}$ thermodynamic degrees of freedom for the black hole. A single Planckon has no partner to form a pair with, and so has zero entropy:
$S(N=1)=S({\cal N}=0)=0\,.$ (8)
That it is the pairs, and not the Planckons themselves, that are the black hole’s thermodynamic degrees of freedom can be seen from the equipartition law, which holds for large ${\cal N}$:
$M=2{\cal N}T_{\rm H}\,\,,\,\,{\cal N}\gg 1\,.$ (9)
Here $T_{\rm H}$ is the Hawking temperature:
$T_{\rm H}=\frac{1}{8\pi GM}=\frac{m_{\rm P}}{N}\,.$ (10)
II.3 Splitting of black hole into smaller parts and back reaction
The discrete version of the Tsallis-Cirto $\delta=2$ entropy determines the processes of merging and splitting of ensembles of Planckons. For example, the probability of a black hole splitting into two smaller parts, $N=N_{1}+N_{2}$, in the process of macroscopic quantum tunneling is determined by the decrease in entropy after the split:
$w(N\rightarrow N_{1}+N_{2})=\exp[S_{\rm BH}(N_{1})+S_{\rm BH}(N_{2})-S_{\rm BH}(N_{1}+N_{2})]=\exp(-N_{1}N_{2})\,.$ (11)
In particular, the rate of emission of one Plankcon by a black hole is:
$w(N\rightarrow(N-1)+1)=e^{-(N-1)}\,.$ (12)
It can be written in the form corresponding to thermal radiation of Planckons with Hawking temperature:
$w=e^{-(N-1)}=\exp\left(-\frac{m_{\rm P}}{T_{\rm H}}\left(1-\frac{m_{\rm P}}{M}\right)\right)\,.$ (13)
The correction to the Hawking radiation rate corresponds to the effect of back reaction in the process of radiation. This effect was first considered by Parikh and Wilczek,Wilczek2000 the extension of their theory to the processes of splitting of the black holes see in Section II.F of Ref. Volovik2022 .
Similarly, we can consider the processes of division of a black hole into more than two parts. The probability of division into four smaller parts contains $4\times(4-1)/2=6$ exponents:
$w(N\rightarrow N_{1}+N_{2}+N_{3}+N_{4})=\exp(-N_{1}N_{2}-N_{1}N_{3}-N_{1}N_{4}-N_{2}N_{3}-N_{2}N_{4}-N_{3}N_{4})\,.$ (14)
Then the splitting of a black hole into $k$ black holes contains $k/(k-1)/2$ exponents, and finally, the probability of splitting of the black hole into its $N$ Planckons contains ${\cal N}=N(N-1)/2$ exponents:
$w(N\rightarrow\underbrace{1+1+...+1}_{N}\,)=e^{-\cal N}=\exp\left(-\frac{N(N-1)}{2}\right)=\exp(-S_{\rm BH}(N))\,.$ (15)
According to Eq.(12), this process can be considered as the sequence of radiation of Planckons one by one until the complete destruction of the black hole:
$w(N\rightarrow\underbrace{1+1+...+1}_{N}\,)=\Pi_{k=1}^{N}e^{-(N-k)}=\exp\left(-\frac{N(N-1)}{2}\right)=\exp(-S_{\rm BH}(N))\,.$ (16)
In this process the entropy decreases from $S_{\rm BH}(N)=N(N-1)/2$ to $S=0$.
Equations (15) and (16) give the following physical meaning of the entropy of a black hole, regardless of the model. The entropy of a black hole determines the probability of the complete annihilation of the black hole.
III White holes
III.1 Probability of formation of white hole and negative entropy
The white hole has extremely fine-tuned nature as was discussed by Witten.Witten2025 ; Eardley1974 The formation of the white hole from $N$ Planckons is extremely unlikely. The white hole metric is obtained from the metric of the black hole by the time reversal, at which the shift vector ${\bf v}$ of PG coordinates changes sign, $T{\bf v}=-{\bf v}$. Due to this time anti-symmetry between the black and white holes, the probability of formation of the white hole by combining $N$ Planckons is the same as probability of black hole splitting to $N$ Planckons in Eqs. (15) and (16):
$w(\,\underbrace{1+1+...+1}_{N}\,\rightarrow N)=\exp\left(-\frac{N(N-1)}{2}\right)=e^{-\cal N}\,.$ (17)
This means that formation of the white hole by merging of Planckons represents the rare thermodynamic fluctuation, at which the entropy decreases from $S=0$ to the maximum negative value, which corresponds to the entropy of a black hole with a minus sign:
$S_{\rm WH}(N)=-\frac{N(N-1)}{2}=-{\cal N}=-S_{\rm BH}(N)\,\,,\,\,N=\frac{M}{m_{\rm P}}\,.$ (18)
This equation suggests the analog of Bekenstein entropy boundBekenstein1981 ; Bekenstein2004 and holographic entropy boundHooft1993 ; Susskind1995 , which now includes the negative entropy of white holes. The entropy of an ensemble of $N$ Planckons is bounded by maximum and minimum values that correspond to the entropies of black and white holes:
$-\frac{N(N-1)}{2}\leq S(N)\leq\frac{N(N-1)}{2}\,.$ (19)
The entropy of a white hole is the negative entropy of the least probable exceptional state. A white hole relaxes into a black hole with increasing entropy, and thus due to the negative entropy of a white hole, the second law of thermodynamics is not violated.
Eq. (18) is supported by the direct calculations of the transition from the black hole to the white hole by macroscopic quantum tunneling,Volovik2022 which demonstrate that the entropy of the white hole with mass $M$ is with minus sign the entropy of the black hole with the same mass:
$S_{\rm WH}(M)=-S_{\rm BH}(M)=-4\pi GM^{2}\,.$ (20)
In the same manner the time reversal transforms the expanding de Sitter with the positive Hubble parameter, $H>0$, to contracting de Sitter with $H<0$. Correspondingly, due to anti-symmetry of the Hubble parameter, $TH=-H$, the entropy of the volume inside the cosmological horizon (the Hubble volume) is negative in the contracting Universe, $S(H<0)=-S(H>0)=-A/4G$, where $A$ is the area of the cosmological horizon.Volovik2025b In transition between expanding and contracting Universe via Minkowskii state, the local entropy density $s$ changes sign by crossing zero value, which is natural for the empty Minkowski vacuum. But the total entropy $S(H)=sV_{H}$ of the Hubble volume $V_{H}$ changes sign by crossing an infinite value. This shows that spatial infinity in the Minkowski vacuum can play a role similar to that of the cosmological event horizon. Escape to infinity may play the same role as escape through the crossing of the cosmological horizon.
III.2 Tsallis-Cirto statistics with negative entropy and second law of thermodynamics
In all the processes of splitting of the black hole, the entropy decreases, as it does in rare thermodynamic fluctuations.Landau_Lifshitz The process of macroscopic quantum tunneling, in which a black hole transforms into a white hole, can also be viewed as a thermodynamic fluctuation in which the entropy decreases. Thermodynamic fluctuations are rare events, they do not violate the second law of thermodynamics since this negative jump of the entropy after splitting is compensated by the further processes of absorption of the radiated objects which restore the entropy of the black hole.
On the other hand, the natural process of splitting of the white hole into smaller parts is always accompanied by increase of entropy. The increase of entropy does not require fluctuations in the form of macroscopic quantum tunneling. That is why the white hole also does no violate the second law of thermodynamics.
So, the black and white holes are thermodynamic objects of special type. Due to the presence of horizon, they obey the special type of the thermodynamic statistics – the extended Tsallis-Cirto $\delta=2$ statistics with
$S^{\pm}_{\delta=2}=\pm\sum_{i}p_{i}\left(\ln\frac{1}{p_{i}}\right)^{2}\,.$ (21)
This differs from both the usual von Neumann statistics, and also from the original $\delta=2$ Tsallis-Cirto statistics in Eq.(4)) by allowing a negative value for the entropy of the white-hole horizon.Volovik2025d States with negative thermodynamic entropy are forbidden in ordinary thermodynamics of systems without horizons, but are natural for thermodynamics of systems with horizons.
IV Reissner-Nordström black hole
IV.1 Negative entropy of inner horizon and composition law
In the extended Thallis-Cirto statistics in Eq.(21), the entropy of Reissner-Nordström (RN) black hole can be considered as the composition of the entropies of the outer and inner horizons:
$S_{\rm RN}=\left(\sqrt{S_{\rm RN}(r_{+})}+\sqrt{|S_{\rm RN}(r_{-})|}\right)^{2}=4\pi GM^{2}\,.$ (22)
Here entropy of the outer horizon is
$S_{\rm RN}(r_{+})=\pi r_{+}^{2}/G=\pi G\left(M+\sqrt{M^{2}-\alpha Q^{2}}\right)^{2}\,,$ (23)
where $\alpha$ is the fine structure constant, and $Q$ is the integer valued electric charge with $Q=-1$ for electron, and the entropy of the inner horizon is negative:
$S_{\rm RN}(r_{-})=-\pi r_{-}^{2}/G=-\pi G\left(M-\sqrt{M^{2}-\alpha Q^{2}}\right)^{2}\,.$ (24)
The minus sign for the entropy of the inner horizon in Eq. (24) is due to the fact that the inner horizon is equivalent to the white hole horizon and therefore its entropy is negative. The negative entropy of the inner horizon also agrees with the negative temperature of the inner horizon, $T_{-}=-(r_{+}-r_{-})/4\pi r_{-}^{2}$, see Eq.(53) in Ref. Volovik2022 .
The composition rule for the Reissner-Nordström black hole shows that the total entropy of an RN black hole is completely determined by its mass $M$. This is consistent with the adiabatic transformation of the RN black hole into a Schwarzschild black hole, where the parameter $\alpha$ (the fine structure constant) adiabatically transforms to zero for fixed $M$ and $Q$. During this adiabatic transformation, the entropy does not change.
In the Planckon description, the RN black hole can be considered as ensemble of of Planckons, each with charge $q=Q/N$ (the charged Planckons have been also considered by Markov.Markov1967 ). Then Eq.(22) demonstrates that the entropy of the RN ensemble does not depend on the Planckon charge. The entropy of RN black hole depends only on the number ${\cal N}$ of degrees of freedom, and thus on the number $N$ of Planckons.
$S_{\rm RN}={\cal N}=\frac{N(N-1)}{2}\,.$ (25)
IV.2 Extremality and fine structure constant
Equation (25) is valid if the RN black hole is non-extremal. The transition to a super-critical black hole by continuously changing the fine structure constant $\alpha$ is non-adiabatic, as can be seen from the entropy $S_{+}$ and $S_{-}$ of outer and inner horizons:
$S_{\pm}=\pm\frac{N(N-1)}{8}\left(1\pm\sqrt{1-8\pi\alpha q^{2}}\right)^{2}\,\,,\,\,q=\frac{Q}{N}\,.$ (26)
The continuity is violated at the critical value, $\alpha_{c}=1/(8\pi q^{2})$. This discontinuity can be explained in terms of gravitational and Coulomb interactions between Planckons:
$U(r)=\frac{\alpha q^{2}}{r}-\frac{Gm_{\rm P}^{2}}{r}\,.$ (27)
At a critical value $\alpha=\alpha_{c}$ the interaction disappears, $U(r,\alpha_{c})=0$. At larger $\alpha>\alpha_{c}$ the interaction between Planckons becomes repulsive, which leads to instability. The Planckons fly apart, emptying the black hole, and the entropy drops to zero.
At fixed $\alpha$ the transition to super-criticality is determined by the critical charge of Planckon:
$q_{c}=\frac{1}{\sqrt{8\pi\alpha}}\,.$ (28)
For the current value of the fine structure constant, the critical charge is about two electron charges, $q_{c}\sim 2$.
IV.3 Horizons, thermodynamics and Lifshitz transition
So, the approach to extremality is not continuous,Carroll2009 and the entropy of the RN black hole drops from ${\cal N}=N(N-1)/2$ to zero when two horizons merge and annihilate. In the same manner the entropy of the RN white hole jumps from its negative value $-{\cal N}=-N(N-1)/2$ to zero when horizons annihilate. This shows that the negative entropy only holds in the presence of horizons.
At the critical point, at $\alpha_{c}=1/8\pi q^{2}$, adiabatic principle is violated. The discontinuous transition between two regimes can be considered as a type of topological quantum phase transitionsVolovik2007 in which horizons are formed, destroyed, annihilate, or change their topology.Jacobson1995 ; Smolin2003 Here the horizon represents the space-time analogue of the Fermi surface, which undergoes Lifshitz transitions.Lifshitz1960
The event horizon is the source of irreversibility, which is similar to the statistical irreversibility.Witten2025 The emergence of horizons corresponds to the spontaneous breaking of time reversal symmetry, which leads to irreversibility. Without horizons, PG metrics with opposite shift vectors ${\bf v}$ describe the same state, since the two metrics can be transformed into each other by continuous coordinate transformations. This is why the metric becomes symmetric under time reversal and thus there is no irreversibility. It is the event horizons that provide positive temperature and entropy for a black hole and negative temperature and entropy for a white hole. This can be seen from the specific detailed balance in the presence of horizon:
$P_{\rm emission}=P_{\rm absorption}\exp(-\omega/T)\,,$ (29)
where for simplicity we consider the limit $\omega\gg T$.
Here by absorption we mean the absorption of external matter by the black hole, which is equivalent to conventional spontanous emission from an excited state. In the presence of a black hole horizon, matter is inevitably absorbed with probability $P_{\rm absorption}=1$, and emission occurs only through quantum tunneling from the black hole with probability $P_{\rm emission}=\exp(-\omega/T_{\rm Hawking})$. Then the detailed balance (29) gives the Hawking temperature for black hole, $T=T_{\rm Hawking}=1/8\pi M$. For the white hole horizon the emission is inevitable, $P_{\rm emission}=1$, while now absorption is provided by quantum tunneling to white hole, $P_{\rm absorption}=\exp(-\omega/T_{\rm Hawking})$. This corresponds to negative temperature $T=-T_{\rm Hawking}=-1/8\pi M$.
In the case of a RN black hole we have $P_{\rm absorption}=1$ and $P_{\rm emission}=\exp(\omega/|T_{-}|)\exp(-\omega/T_{+})$. Then the detailed balance (29) gives $1/T=1/T_{+}-1/|T_{-}|=4\pi(r_{+}+r_{-})=8\pi M$. This again supports the conclusion that the thermodynamics of the RN black hole coincides with the thermodynamics of the electrically neutral black hole with the same mass.
V de Sitter thermodynamics in Planckon model and quantization of $\Lambda$
The equation (29) is also applicable to the processes associated with the cosmological horizon. The corresponding temperature governing the processes of radiation from the horizon is the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$, where $H$ is the Hubble parameter. The local temperature of the de Sitter state, describing the de Sitter thermodynamics, is not related to horizon and is twice as large.Volovik2025b Nevertheless, the local entropy density $s$ integrated over the Hubble volume $V_{H}$ (the volume inside the cosmological horizon) corresponds to the Gibbons-Hawking entropy of the cosmological horizon, $S_{H}=sV_{H}=A/4G$, where $A$ is the area of the cosmological horizon. This connection between local thermodynamics and the thermodynamics of the cosmological horizon is a consequence of the specific symmetry of the de Sitter state and is supported by topological arguments.Kusmartsev2025 ; Volovik2025f
Another connection is provided by the Planckon model. According to this model, the energy and entropy of the Hubble volume are quantized in the same way as in the case of a black hole:
$E_{H}=Nm_{\rm P}\,\,,\,\,S_{H}={\cal N}=\frac{N(N-1)}{2}\,.$ (30)
For large $N\gg 1$ the entropy of the Hubble volume approaches the Gibbons-Hawking entropy:
$S_{H}(N\gg 1)\rightarrow\frac{1}{2}N^{2}=\frac{\pi}{GH^{2}}=\frac{A}{4G}\,,$ (31)
On the other hand, quantized entropy implies quantization of the cosmological constant $\Lambda$, i.e. the energy density $\rho_{\rm vac}$ of the de Sitter state, which is a local quantity:
$\Lambda\equiv\rho_{\rm vac}=\frac{3}{8G^{2}{\cal N}}=\frac{24\pi^{2}}{l_{\rm P}^{4}{\cal N}}\,.$ (32)
Here $l_{\rm P}=1/m_{\rm P}$ is the reduced Planck length. A similar quantization of $\Lambda$ is obtained in the gravitational-topological sector.Wieland2011 ; Alexander2025 Equation (23) in Ref.Wieland2011 implies for the reduced cosmological constant the values $\Lambda=\frac{12\pi}{\beta l_{\rm P}^{2}n}$, where $\beta$ is the Barbero–Immirzi parameter and $n$ is an integer. With $\beta=\frac{1}{2\pi}$ this agrees with Eq.(32).
Let us also mention the quantum theory of gravity, in which tetrads are formed as bilinear combinations of fermion fields and have the dimension of the inverse length.Diakonov2012 As a result, the cosmological constant $\Lambda$ is dimensionless, since $l_{\rm P}^{4}$ is absorbed by the 4-from tetrad determinant. For dimensionless physical quantities, quantization is quite natural. In this theory of gravity, the covariant form of the scalar area is also dimensionless, which makes quantization of the horizon area and thus of the horizon entropy also rather natural.Volovik2021
VI Conclusion
Thermodynamics of black and white holes can be considered in terms of the ensemble of the Planck mass objects, Planckons. The Tsallis-Cirto non-extensive statistics with $\delta=2$, which describes the processes of splitting and merging of black holes, suggests the statistical properties of this ensemble. The entropy of black hole with $N$ Planckons is provided by the $N(N-1)/2$ degrees of freedom – the correlations between the gravitationally attracted Planckons. If one chooses the mass of Planckon as the reduced Planck mass, $m_{\rm P}=1/\sqrt{8\pi G}$, and thus the mass of the black hole is $M=Nm_{\rm P}$, then the entropy of this black hole is $S_{\rm BH}(N)=N(N-1)/2$.
This describes the processes of splitting and merging of black holes in integer numbers. For example, the probability of splitting of the black hole with $N$ Planckons to the two black holes $N_{1}+N_{2}=N$ is $w\propto e^{-N_{1}N_{2}}$. For $N_{1}=1$ and $N_{2}=N-1$ this gives the probability of thermal radiation of Planckon with Hawking temperature $T_{\rm H}=1/8\pi GM$ supplemented by the back-reaction effect in the radiation process, which first considered by Parikh and Wilczek. The probability of the decay of a black into its $N$ Planckons is $w\propto e^{-N(N-1)/2}$, which relates the entropy of a black hole to the probability of its complete annihilation.
The entropy of a white hole is negative as follows from the rate of macroscopic quantum tunneling from a black hole to a white hole. In terms of Planckons, the entropy of a white hole is the negative of the entropy of a black hole with the same number of Planckons, $S_{\rm WH}(N)=-N(N-1)/2=-S_{\rm BH}(N)$. This reflects the anti-symmetry with respect to time reversal, at which the shift vector in the Arnowitt-Deser-Misner formalism changes sign. This anti-symmetry allows one to extend the Tsallis-Chirto entropy by adding a minus sign to the Tsallis-Chirto formula applied to white hole.
The Planckon model has been extended to a Reissner-Nordström (RN) black hole, which consists of the electrically charged Planckons. The entropy of a Reissner-Nordström black hole is obtained as a non-extensive composition of the positive entropy of the outer horizon and the negative entropy of the inner horizon. Despite the charge, the statistical ensemble of Planckons remains the same, and the RN black hole with $N$ Planckons has the same entropy as the electrically neutral hole, $S_{\rm RN}(N)=N(N-1)/2$. This is confirmed by the adiabatic process of transition from RN to a Schwarzschild black hole by continuous changing the fine structure constant. Adiabaticity is broken in the extreme limit, when the gravitational interaction between two Planckons is compensated by the repulsion between their electric charges, and the RN black hole loses stability and disintegrates into Planckons.
The Planckon model, when extended to the de Sitter state of the quantum vacuum, implies quantization of the cosmological constant $\Lambda$. This connects the thermodynamics of the cosmological horizon with the local thermodynamics of the de Sitter state, as well as with the topological sector of quantum gravity.
According to Unruh,Unruh2024 the quantum mechanics of black holes is still a far from finished topic. Also according to Jacobson,Jacobson2024 the connection between the de Sitter horizon entropy and quantum field theory still stands on shaky ground. This toy model will likely give us a deeper understanding of the origins of thermodynamics that arise in black holes, as well as in the quantum vacuum.